# A probe of the maximum energetics of fast radio bursts through a prolific repeating source

### v1

@@ Abstract
Fast radio bursts (FRBs) are sufficiently energetic to be detectable from luminosity distances up to at least seven billion parsecs (redshift $z>1$). Probing the maximum energies and luminosities of FRBs constrains their emission mechanism and cosmological population. Here we investigate the maximum energetics of a highly active repeater, FRB 20220912A, using 1,500 h of observations. We detect $130$ high-energy bursts and find a break in the burst energy distribution, with a flattening of the power-law slope at higher energy. This is consistent with the behaviour of another highly active repeater, FRB 20201124A. Furthermore, we model the rate of the highest-energy bursts and find a turnover at a characteristic spectral energy density of $E^{\textrm{char}}_{\nu}=2.09^{+3.78}_{-1.04}\times 10^{32}$ erg/Hz. This characteristic maximum energy agrees well with observations of apparently one-off FRBs, suggesting a common physical mechanism for their emission. The extreme burst energies push radiation and source models to their limit.
Introduction
@@
Fast radio bursts1 (FRBs) are observed with durations ranging from microseconds2 to seconds3 and fluences ranging from about 0.01 Jy ms to 1000 Jy ms4, 5. Of the thousands of known FRBs, most are observed as one-off events and only a few percent are known to repeat6. It remains unclear whether the repeaters and apparent non-repeaters have different astrophysical origins. An exceptionally bright (MJy) burst has been observed from the Galactic magnetar SGR 1935+2154, and strongly suggests that at least some FRBs originate from magnetars7, 8.
@@
Though FRB emission is almost certainly beamed, the beaming fraction is unknown. For simplicity, isotropic-equivalent energies are used to compare different FRBs even though this is an overestimation of the total energy released in radio waves. Given the typical Mpc- to Gpc-distances of FRBs9, their isotropic-equivalent energies are known to range from roughly $E=10^{37}-10^{42}$ erg. This is at least four orders-of-magnitude larger than what is measured for Galactic radio pulsars, including the giant pulses seen from the Crab pulsar10. SGR 1935+2154, however, is known to produce bursts whose energies span the range between pulsars and FRBs (Ref. 11 and references therein).
@@
Furthermore, it is expected that the detected radio energy is only a small fraction ($\sim 10^{-5}$) of the total bolometric energy released at the source during an FRB event12, 13. Taking this into account, the most energetic known FRBs may be associated with $\sim 10^{46-47}$ erg events that are comparable to the most extreme ‘giant flares’ from magnetars14, though still at least four orders-of-magnitude less energetic compared to supernovae and gamma-ray bursts. Nonetheless, FRBs display extreme brightness temperatures $T_{\rm B}\gg 10^{12}$ K (the typically assumed threshold between incoherent and coherent emitters), and therefore must originate from a coherent emission process2. This makes them detectable across cosmic distances despite their relatively modest energetics compared to other extreme astrophysical phenomena.
@@
The distribution of observed FRB spectral energies ($E_{\nu}$) follow a differential power law, where the rate above some spectral energy scales as $R~{}(>E_{\nu})\propto E_{\nu}^{\gamma_{D}}$ for a differential index $\gamma_{D}$. For the overall population of apparently non-repeating sources15, 16, $\gamma_{D}\sim-1.5$. Highly active repeating FRBs provide the opportunity to measure the burst energy distribution of a single source. In the case of repeaters, the energy distribution has been shown to deviate from a simple power law17, 5.
@@
The most energetic FRBs are also by far the most rare. Therefore, large on-sky time is essential to probe the extremes of the FRB population. By probing the maximum energetics of FRBs, we can constrain the emission mechanism18, 19; the total cosmic population20, 15, 16; and inform how best to detect FRBs with upcoming telescopes.
@@
There are various conceivable ways to investigate the maximum achievable energies and luminosities of FRBs, including 1. discovery of exceptionally distant or energetic one-off bursts21; 2. population modelling of the full observed sample of FRBs15, 16; and 3. high-cadence monitoring of hyperactive repeating sources5, 22 (those repeaters that are sometimes seen to produce hundreds of bursts per hour, if observed with a high-sensitivity radio telescope). These methods complement each other, given that they are all subject to different observational biases and challenges. Moreover, it remains unclear whether apparently one-off and repeating FRBs share the same progenitors and emission mechanisms. By ‘progenitor’, we mean the type of astrophysical source powering the bursts, and by ‘emission mechanism’ we mean the physical process that generates the bursts. Probing the maximum observed energies of repeaters and apparently one-off FRBs is thus also a way to compare their nature.
[…]
Observations, analysis and results

### v2

@@ Abstract
Fast radio bursts (FRBs) are sufficiently energetic to be detectable from luminosity distances up to at least seven billion parsecs (redshift $z>1$). Probing the maximum energies and luminosities of FRBs constrains their emission mechanism and cosmological population. Here we investigate the maximum energetics of a highly active repeater, FRB 20220912A, using 1,500 h of observations. We detect $130$ high-energy bursts and find a break in the burst energy distribution, with a flattening of the power-law slope at higher energy – consistent with the behaviour of another highly active repeater, FRB 20201124A. There is a roughly equal split of integrated burst energy between the low- and high-energy regimes. Furthermore, we model the rate of the highest-energy bursts and find a turnover at a characteristic spectral energy density of $E^{\textrm{char}}_{\nu}=2.09^{+3.78}_{-1.04}\times 10^{32}$ erg Hz-1. This characteristic maximum energy agrees well with observations of apparently one-off FRBs, suggesting a common physical mechanism for their emission. The extreme burst energies push radiation and source models to their limit: at this burst rate a typical magnetar ($B=10^{15}$ G) would deplete the energy stored in its magnetosphere in $\sim$ 2150 h, assuming a radio efficiency $\epsilon_{\mathrm{radio}}=10^{-5}$. We find that the high-energy bursts ($E_{\nu}>3\times 10^{30}$ erg Hz-1) play an important role in exhausting the energy budget of the source.
keywords:
fast radio bursts – radio continuum: transients
pubyear: 2025pagerange: A probe of the maximum energetics of fast radio bursts through a prolific repeating sourceB.2
1 Introduction
@@
Fast radio bursts (FRBs; Petroff et al., 2022) are observed with durations ranging from microseconds (Snelders et al., 2023) to seconds (CHIME/FRB Collaboration et al., 2022) and fluences ranging from about 0.01 Jy ms to 1000 Jy ms (Spitler et al., 2016; Kirsten et al., 2024). Of the thousands of known FRBs, most are observed as one-off events and only a few percent are known to repeat (CHIME/FRB Collaboration et al., 2023). It remains unclear whether the repeaters and apparent non-repeaters have different astrophysical origins. An exceptionally bright (MJy) burst has been observed from the Galactic magnetar SGR 1935+2154, and strongly suggests that at least some FRBs originate from magnetars (CHIME/FRB Collaboration et al., 2020; Bochenek et al., 2020).
@@
Though FRB emission is almost certainly beamed, the beaming fraction is unknown. For simplicity, isotropic-equivalent energies are used to compare different FRBs even though this is an overestimation of the total energy released in radio waves. Given the typical Mpc- to Gpc-distances of FRBs (Gordon et al., 2023), their isotropic-equivalent energies are known to range from roughly $E=10^{37}-10^{42}$ erg. This is at least four orders-of-magnitude larger than what is measured for Galactic radio pulsars, including the giant pulses seen from the Crab pulsar (Bera & Chengalur, 2019). SGR 1935+2154, however, is known to produce bursts whose energies span the range between pulsars and FRBs (Ref.  (Kirsten et al., 2021) and references therein).
@@
Furthermore, it is expected that the detected radio energy is only a small fraction ($\sim 10^{-5}$) of the total bolometric energy released at the source during an FRB event (Metzger et al., 2019; Mereghetti et al., 2020). Taking this into account, the most energetic known FRBs may be associated with $\sim 10^{46-47}$ erg events that are comparable to the most extreme ‘giant flares’ from magnetars (Kaspi & Beloborodov, 2017), though still at least four orders-of-magnitude less energetic compared to supernovae and gamma-ray bursts. Nonetheless, FRBs display extreme brightness temperatures $T_{\rm B}\gg 10^{12}$ K (the typically assumed threshold between incoherent and coherent emitters), and therefore must originate from a coherent emission process (Snelders et al., 2023). This makes them detectable across cosmic distances despite their relatively modest energetics compared to other extreme astrophysical phenomena.
@@
The distribution of observed FRB spectral energies ($E_{\nu}$) follow a differential power law, where the rate above some spectral energy scales as $R~(>E_{\nu})\propto E_{\nu}^{\gamma_{D}}$ for a differential index $\gamma_{D}$. For the overall population of apparently non-repeating sources (James et al., 2022b; Shin et al., 2023), $\gamma_{D}\sim-1.5$. Highly active repeating FRBs provide the opportunity to measure the burst energy distribution of a single source. In the case of repeaters, the energy distribution has been shown to deviate from a simple power law (Li et al., 2021b; Kirsten et al., 2024).
@@
The most energetic FRBs are also by far the most rare. Therefore, large on-sky time is essential to probe the extremes of the FRB population. By probing the maximum energetics of FRBs, we can constrain the emission mechanism (Lu et al., 2019; Cooper & Wijers, 2021); the total cosmic population (Luo et al., 2020; James et al., 2022b; Shin et al., 2023); and inform how best to detect FRBs with upcoming telescopes.
@@
There are various conceivable ways to investigate the maximum achievable energies and luminosities of FRBs, including 1. discovery of exceptionally distant or energetic one-off bursts (Ryder et al., 2023); 2. population modelling of the full observed sample of FRBs (James et al., 2022b; Shin et al., 2023); and 3. high-cadence monitoring of hyperactive repeating sources (Kirsten et al., 2024; Sheikh et al., 2024) (those repeaters that are sometimes seen to produce hundreds of bursts per hour, if observed with a high-sensitivity radio telescope). These methods complement each other, given that they are all subject to different observational biases and challenges. Moreover, it remains unclear whether apparently one-off and repeating FRBs share the same progenitors and emission mechanisms. By ‘progenitor’, we mean the type of astrophysical source powering the bursts, and by ‘emission mechanism’ we mean the physical process that generates the bursts. Probing the maximum observed energies of repeaters and apparently one-off FRBs is thus also a way to compare their nature.
In this work we investigate the maximum burst energy and luminosity of FRB 20220912A, using multiple 25-32 m–class radio telescopes that together provide unprecedented observational coverage in terms of on-sky time. FRB 20220912A was discovered using the CHIME/FRB system (Mckinven & CHIME/FRB Collaboration, 2022), and was soon identified as a hyperactive source compared to most other known repeaters. It was localised to a host galaxy at $z=0.0771$ (Ravi et al., 2023; Hewitt et al., 2024) and has been the target of many follow-up observations. The exceptionally high activity of FRB 20220912A makes it an ideal source to map the burst energy distribution of a repeater. Here we specifically focus on the maximum achievable burst energy and luminosity.
2 Observations